OpenStudy (anonymous):

Solve: −2(2x + 5) − 3 = −3(x − 1) help?

OpenStudy (anonymous):

First distribute: -4x-10-3 = -3x+3 Then solve: -4x -13 = -3x+3 -x-13=3 -16 = x

hero (hero):

−2(2x + 5) − 3 = −3(x − 1) -2(x + x - 1 + 6) - 3 = -3(x - 1) -2x -2(x - 1) - 12 - 3 = -3(x - 1) -2x + 3(x - 1) - 2(x - 1) - 15 = 0 -2x + x - 1 = 15 -x = 15 + 1 -x = 16 x = -16

OpenStudy (anonymous):

thxs guys! :)

hero (hero):

My steps are a bit difficult to digest.